Please help! I cannot figure out how to set this problem up. I have to include the cost function as well. Thank you! Here it is:
An appliance repair service charges $50 to come out for a service call which is applied toward the first 30 minutes of service should a repair be needed. If a repair is needed, the charge of $37 per 30 minutes (after the initial half hour included in the service call charge) for labor and the cost of parts. Write a cost function to determine the amount a customer needing a repair that takes 1.5 hours and requires a $103 part.

Answer by checkley71(8403)

50+(2*37)+103=50+74+103=$227 FOR THE TOTAL SERVICE BILL.